OpenStudy (anonymous):

1.) 2x + y = 3 2.) x - 2y = -1 If equation 1 is multiplied by 2 and then the equations are added, the result is

OpenStudy (astrophysics):

\[(2x+y=3)*2 =...\] then add it to the second equation